OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there? HELP!!!!!!!!

OpenStudy (ash2326):

Let the no. of humans be x and no. of horses be y so We have 74 head altogether \[x+y=74\] Do you agree with this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

How many legs does a human have? What about horse?

OpenStudy (anonymous):

Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 So, the solution is 24 horses and 50 humans.

OpenStudy (anonymous):

thanks dude

OpenStudy (ash2326):

cool :)